# Evaluating a VR System for Collecting Safety-Critical Vehicle-Pedestrian Interactions

### v1

@@ 5.3.1. Quantitative Measures
We report the average user ratings for each question in Table 3.
@@
Table 3. VR Presense Questionnaire Evaluation Results.
5.3.2. Qualitative Measures
@@
The qualitative responses from our user study affirm that, while the system does have areas for improvement, there are noteworthy strengths. Participants largely found the visual aspects and their own emotional responses to be authentic to their real-world experience. Some users did report aspects of the virtual environment that seemed unrealistic; yet sense of presence, agency, and behavioral similarity to real-life still seems high (Table 3). This seems to suggest that complete realism of the virtual environment may not be necessary to have high sense of presence,
@@
6. Part 3: Third-Person Realism Evaluation of VR vs. Real-Life vs. Synthetic 2D Trajectory Data
@@
6.1. Study Design
To collect real-life jaywalking trajectories, we visited a two-way single-lane street near the university that is well-trafficked by jaywalkers as well as vehicles. We chose this setting for its potential to provide real-life trajectories in a similar geographic layout to that we used in the VR jaywalking scenario. In contrast to the VR scenario, there were fewer vehicles as well as more irregular gaps between vehicles, making it easier for jaywalkers to find a gap in which to cross. A DJI Mavic Mini drone (noa, [n. d.]) was used for data collection, flown at a height of approximately 25 meters. In a 1-hour time interval, we secured footage for 7 jaywalkers.
@@ 6.1. Study Design
The trajectories were rendered as animated gif images, in which the pedestrian was depicted as a small red circle, and vehicles as blue rectangles. A random rotation was added to each image to eliminate visual bias due to the layout of the scene. The spatial limits of the animation were set such that the users had an ample field of view to see oncoming vehicles from both sides (which is important for judging pedestrian behavior). All animations were set to have a similar field of view. For the real-life trajectories, only the single jaywalking pedestrian was annotated. Scenes were chosen such that the pedestrians kept an ample distance from other pedestrians such that they did not visibly interact or influence each other’s motion; i.e., the vehicle-pedestrian interaction would be the sole interaction evaluated. An example schematic in the style of the animations is shown in Figure 6.
Figure 6. A frame from an example schematic shown to survey respondents.
@@ 6.2. Results and Analyses
Consistent with expectations about the “control” variable, evaluators could easily tell the difference between a simple synthetic policy and a genuine human trajectory. 89.3% of real-life / synthetic comparisons were evaluated in favor of the real-life trajectory looking “more real” than the synthetic. The response was similarly high for the VR / synthetic comparisons, with 88.9% evaluated in favor of the virtual over the synthetic (Table LABEL:tab:3rd_person).
Also consistent with expectations, the real-life trajectories were the most-frequently chosen as “more real,” with approximately 67% of responses choosing it as more “real” in a pair (Table LABEL:tab:3rd_person). However, VR trajectories do not lag far behind; 59% of responses chose the VR trajectories as “more real” in a pair (Table LABEL:tab:3rd_person). Of particular note is that 36.5% of real-life / VR comparisons were evaluated in favor of the VR trajectory looking “more real” (Table 5); these responses claimed that the VR trajectory looked “more real” than a real trajectory.
Though 36.5% is still a bit away from 50%, which would mean that real-life trajectories are indistinguishable from VR trajectories, this number still substantiates the claim that VR trajectories can evoke genuine pedestrian responses. First of all, the results confirm that real-life trajectories are less distinguishable from VR trajectories than they are from fully-synthetic trajectories, supporting the claim that the VR system evokes more natural pedestrian responses than simple policies used by synthetic pedestrian simulators (Biswas et al., 2021; Kothari et al., 2022). Second, there could be other reasons evaluators are able to tell the difference between real and VR trajectories that have nothing to do with how ”real” VR trajectories are. For example, one respondent who evaluated 3 comparisons between real-life and VR trajectories and marked all the real-life ones as more “real” gave the free response answer that “the fake pedestrians seem to cross dangerously close to the vehicles, and sometimes stop in the middle of the road…” Though this respondent labelled all real-life trajectories as “more real” than the corresponding VR ones, their explanation for their rating is not necessarily true. It is generally considered dangerous and not commonly observed that pedestrians “cross dangerously close to the vehicles.” As observed in all drone-collected videos, the average jaywalker in real-life is a “safe” jaywalker who only jaywalks when there is sufficient gap in the vehicles. However, sometimes there are pedestrians who do engage in risk-taking behavior. Sometimes there are jaywalkers who jaywalk even on heavily-trafficked roads. These jaywalkers, not able to find a gap between vehicles, must stop in the middle of the road before they can cross completely. Though rare, “dangerous” jaywalkers do exist in real-life, and it is important that safety-critical AV systems have their data.
Truth Value $\rightarrow$
Guessed value $\downarrow$
Real-life VR Synthetic
@@
Truth Value $\rightarrow$
Guessed value $\downarrow$
Real-life VR Synthetic
@@
Table 4. Confusion matrix for survey respondent guesses (real / fake) vs. truth category (real-life, VR, synthetic)
@@
Percent Correct (N) 64.5% (732) 89.3% (84) 88.9% (90)
Table 5. Percent correct of each pairing. “Correct” is defined as selecting real as more “real” when compared with either VR or synthetic, and selecting VR as more “real” when compared to synthetic.
7. Limitations, Discussion, and Future Work
7.1. Interview Responses from Part 1 Not Addressed by VR System
@@
  1. (1)
    I1 and I2 suggested that the data from vehicle simulators may be less noisy than real-life data, enlarging the sim-to-real gap. In the VR CARLA environment, vehicles follow simple policies like the Intelligent Driver Model, always driving perfectly along predefined paths. This may not realistically reproduce human driving behavior, which is imperfect, and sometimes deviates from the center of the lane. In the future, more complex and noisy driving policies can be implemented into the CARLA environment to create more realistic vehicle behavior. In addition, as described in the concurrent work contributing the VR system (Anonymized, 2023), a real human driver can be hooked into the simulator as an additional vehicle agent. The real human driver will create vehicle trajectories that pedestrians users may deem more natural.
  2. (2)
@@
7.2. Free Responses from Part 2 that Elicit VR System Limitations
The qualitative responses summarized in Section 5.3.2 to the free response question, What aspects of the system or environment were unrealistic?, revealed certain limitations of the VR system that can be expanded in future work. While some concerns are difficult to address with the state of current technologies, such as (1), (2), and (4), other concerns can be resolved with additional engineering of the VR system or revised study design. To address (3), giving participants a longer time to familiarize themselves with the VR headset may increase their reassurance that they will not run into obstacles in the real world. To address (5), sound can be easily added to the simulation as it is available in CARLA. However, we chose not to use it in our study, so that participants could hear us when we verbally instructed participants to where they should walk. (6) can be addressed by adding into the scene the relevant environmental elements, all of which are available as assets in Unreal Engine. (7) may be a bit more difficult to address across all environments, but can be addressed for the sidewalk / curb situation by adding physical objects in the real world to match those in the virtual world, and carefully calibrating the virtual-to-real mapping. (8) can be addressed by integrating into the system the driving simulator extension, which has already been developed. A human driver participant can then use the driving simulator to operate the vehicles in more natural way. Furthermore, the vehicles currently move based on default CARLA policies; better policies can be integrated into CARLA for more realistic vehicle movement. (10) can be addressed by adding in additional pedestrians into the simulation, which the system also supports; each pedestrian would require an additional VR setup, including headset, wireless tracker, and computer. Alternatively, autonomous pedestrians could be added into the simulator. Although these pedestrians would move based on simple policies, these policies are reactive to the user’s movements. To address (11), participants can be given tasks to do rather than specific goals to navigate to, such as “Please drop off a letter in the nearest mail box,” rather than “Please walk across the road.”
@@ 7.2. Free Responses from Part 2 that Elicit VR System Limitations
7.3. Part 3 Study Design Choice Limitations
@@
8. Conclusion
We evaluated the degree to which a VR data collection system can elicit genuine pedestrian responses, confirming that such a system addresses many of the current limitations of data collection methods. Many aspects of the virtual environment of the system felt real to our users. Further, trajectories collected from the VR system appeared much more realistic than synthetically-generated data to third-person evaluators. Vulnerable road users such as children and elderly behave differently than are Overall, the VR system offers a low-risk method of collecting data of uncommon pedestrian-vehicle interactions, opening the way for vulnerable road users.
9. Conclusion

### v3

@@ V-C1 Quantitative Measures
We report the average user ratings for each question in Table III.
@@
TABLE III: VR Presense Questionnaire Evaluation Results.
V-C2 Qualitative Measures
@@
The qualitative responses from our user study affirm that, while the system does have areas for improvement, there are noteworthy strengths. Participants largely found the visual aspects and their own emotional responses to be authentic to their real-world experience. Some users did report aspects of the virtual environment that seemed unrealistic; yet sense of presence, agency, and behavioral similarity to real-life still seems high (Table III). This seems to suggest that complete realism of the virtual environment may not be necessary to have high sense of presence,
@@
VI Part 3: Third-Person Realism Evaluation of VR vs. Real-Life vs. Synthetic 2D Trajectory Data
@@
VI-A Study Design
To collect real-life jaywalking trajectories, we visited a two-way single-lane street near the university that is well-trafficked by jaywalkers as well as vehicles. We chose this setting for its potential to provide real-life trajectories in a similar geographic layout to that we used in the VR jaywalking scenario. In contrast to the VR scenario, there were fewer vehicles as well as more irregular gaps between vehicles, making it easier for jaywalkers to find a gap in which to cross. A DJI Mavic Mini drone [1] was used for data collection, flown at a height of approximately 25 meters. In a 1-hour time interval, we secured footage for 7 jaywalkers.
@@ VI-A Study Design
Figure 6: A frame from an example schematic shown to survey respondents.
The trajectories were rendered as animated gif images, in which the pedestrian was depicted as a small red circle, and vehicles as blue rectangles. A random rotation was added to each image to eliminate visual bias due to the layout of the scene. The spatial limits of the animation were set such that the users had an ample field of view to see oncoming vehicles from both sides (which is important for judging pedestrian behavior). All animations were set to have a similar field of view. For the real-life trajectories, only the single jaywalking pedestrian was annotated. Scenes were chosen such that the pedestrians kept an ample distance from other pedestrians such that they did not visibly interact or influence each other’s motion; textiti.e., the vehicle-pedestrian interaction would be the sole interaction evaluated. An example schematic in the style of the animations is shown in Figure 6.
@@ VI-B Results and Analyses
Consistent with expectations about the “control” variable, evaluators could easily tell the difference between a simple synthetic policy and a genuine human trajectory. 89.3% of real-life / synthetic comparisons were evaluated in favor of the real-life trajectory looking “more real” than the synthetic. The response was similarly high for the VR / synthetic comparisons, with 88.9% evaluated in favor of the virtual over the synthetic (Table IV).
Also consistent with expectations, the real-life trajectories were the most-frequently chosen as “more real,” with approximately 67% of responses choosing it as more “real” in a pair (Table IV). However, VR trajectories do not lag far behind; 59% of responses chose the VR trajectories as “more real” in a pair (Table IV). Of particular note is that 36.5% of real-life / VR comparisons were evaluated in favor of the VR trajectory looking “more real” (Table V); these responses claimed that the VR trajectory looked “more real” than a real trajectory.
Though 36.5% is still a bit away from 50%, which would mean that real-life trajectories are indistinguishable from VR trajectories, this number still substantiates the claim that VR trajectories can evoke genuine pedestrian responses. First of all, the results confirm that real-life trajectories are less distinguishable from VR trajectories than they are from fully-synthetic trajectories, supporting the claim that the VR system evokes more natural pedestrian responses than simple policies used by synthetic pedestrian simulators [9, 38]. Second, there could be other reasons evaluators are able to tell the difference between real and VR trajectories that have nothing to do with how ”real” VR trajectories are. For example, one respondent who evaluated 3 comparisons between real-life and VR trajectories and marked all the real-life ones as more “real” gave the free response answer that “the fake pedestrians seem to cross dangerously close to the vehicles, and sometimes stop in the middle of the road…” Though this respondent labelled all real-life trajectories as “more real” than the corresponding VR ones, their explanation for their rating is not necessarily true. It is generally considered dangerous and not commonly observed that pedestrians “cross dangerously close to the vehicles.” As observed in all drone-collected videos, the average jaywalker in real-life is a “safe” jaywalker who only jaywalks when there is sufficient gap in the vehicles. However, sometimes there are pedestrians who do engage in risk-taking behavior. Sometimes there are jaywalkers who jaywalk even on heavily-trafficked roads. These jaywalkers, not able to find a gap between vehicles, must stop in the middle of the road before they can cross completely. Though rare, “dangerous” jaywalkers do exist in real-life, and it is important that safety-critical AV systems have their data.
Truth Value $\rightarrow$ Guessed value $\downarrow$ Real-life VR Synthetic
@@
Truth Value $\rightarrow$ Guessed value $\downarrow$ Real-life VR Synthetic
@@
TABLE IV: Confusion matrix for survey respondent guesses (real / fake) vs. truth category (real-life, VR, synthetic)
@@
% Correct (N) 64.5% (732) 89.3% (84) 88.9% (90)
TABLE V: Percent correct of each pairing. “Correct” is defined as selecting real as more “real” when compared with either VR or synthetic, and selecting VR as more “real” when compared to synthetic.
VII Limitations, Discussion, and Future Work
VII-A Interview Responses from Part 1 Not Addressed by VR System
@@
  1. 1.
    I1 and I2 suggested that the data from vehicle simulators may be less noisy than real-life data, enlarging the sim-to-real gap. In the VR CARLA environment, vehicles follow simple policies like the Intelligent Driver Model, always driving perfectly along predefined paths. This may not realistically reproduce human driving behavior, which is imperfect, and sometimes deviates from the center of the lane. In the future, more complex and noisy driving policies can be implemented into the CARLA environment to create more realistic vehicle behavior. In addition, as described in the concurrent work contributing the VR system [3], a real human driver can be hooked into the simulator as an additional vehicle agent. The real human driver will create vehicle trajectories that pedestrians users may deem more natural.
  2. 2.
@@
VII-B Free Responses from Part 2 Eliciting System Limitations
The qualitative responses summarized in Section V-C2 to the free response question, What aspects of the system or environment were unrealistic?, revealed certain limitations of the VR system that can be expanded in future work. While some concerns are difficult to address with the state of current technologies, such as (1), (2), and (4), other concerns can be resolved with additional engineering of the VR system or revised study design. To address (3), giving participants a longer time to familiarize themselves with the VR headset may increase their reassurance that they will not run into obstacles in the real world. To address (5), sound can be easily added to the simulation as it is available in CARLA. However, we chose not to use it in our study, so that participants could hear us when we verbally instructed participants to where they should walk. (6) can be addressed by adding into the scene the relevant environmental elements, all of which are available as assets in Unreal Engine. (7) may be a bit more difficult to address across all environments, but can be addressed for the sidewalk / curb situation by adding physical objects in the real world to match those in the virtual world, and carefully calibrating the virtual-to-real mapping. (8) can be addressed by integrating into the system the driving simulator extension, which has already been developed. A human driver participant can then use the driving simulator to operate the vehicles in more natural way. Furthermore, the vehicles currently move based on default CARLA policies; better policies can be integrated into CARLA for more realistic vehicle movement. (10) can be addressed by adding in additional pedestrians into the simulation, which the system also supports; each pedestrian would require an additional VR setup, including headset, wireless tracker, and computer. Alternatively, autonomous pedestrians could be added into the simulator. Although these pedestrians would move based on simple policies, these policies are reactive to the user’s movements. To address (11), participants can be given tasks to do rather than specific goals to navigate to, such as “Please drop off a letter in the nearest mail box,” rather than “Please walk across the road.”
@@ VII-B Free Responses from Part 2 Eliciting System Limitations
VII-C Part 3 Study Design Choice Limitations
@@
VIII Conclusion